OpenStudy (anonymous):

If V=⟨3,2⟩ and W=⟨-1,-3⟩, find ll V x W ll Can someone please explain how the cross product thing works?

jimthompson5910 (jim_thompson5910):

If Z = <a, b> is the cross product of V and W, then Z is perpendicular to both V and W So this must mean Z dot V = 0 Z dot W = 0

jimthompson5910 (jim_thompson5910):

<a, b> dot <3, 2> = 3a + 2b = 0 <a, b> dot <-1, -3> = -a - 3b = 0 ------------------------------------------------------- So you have these two equations 3a+2b = 0 -a - 3b = 0 which you can use to solve for a and b

jimthompson5910 (jim_thompson5910):

The double vertical bars around V x W tells you to find the magnitude/length of this vector

jimthompson5910 (jim_thompson5910):

oh wait...the cross product is only possible in R3 lol, my bad

OpenStudy (anonymous):

@jim_thompson5910 Cross product only works well in 3 dimensions.

jimthompson5910 (jim_thompson5910):

yeah just remembered lol

jimthompson5910 (jim_thompson5910):

well you can fix that by just adding on a z-coordinate of 1 or 0 or something

OpenStudy (anonymous):

Haha yeah I saw that in my notes and that's why I got confused.. So it is still possible to solve then?

jimthompson5910 (jim_thompson5910):

V = <3, 2> could be thought of as V = <3, 2, 0> W = <-1,-3> could be thought of as W = <-1, -3, 0> since you can think of the xy plane as the plane where z = 0 (in 3 space)

jimthompson5910 (jim_thompson5910):

hmm maybe it might be better to make the third coordinate 1...not sure

OpenStudy (anonymous):

\[ \|\mathbf v \times \mathbf w\| = \|\mathbf v\|\|\mathbf w\|\sin(\theta) \]

OpenStudy (anonymous):

\[ \sin(\theta) = \sqrt{1-\cos^2(\theta)} \]\[ \cos(\theta) = \frac{\|\mathbf v\cdot \mathbf w\|}{\|\mathbf v\|\|\mathbf w\|} \]

jimthompson5910 (jim_thompson5910):

V = <3, 2> could be thought of as V = <3, 2, 1> W = <-1,-3> could be thought of as W = <-1, -3, 1> ------------------------------------------------------- If Z = <a, b, c> then Z dot V = <a, b, c> dot <3, 2, 1> = 3a + 2b + c = 0 Z dot W = <a, b, c> dot <-1, -3, 1> = -a - 3b + c = 0 ------------------------------------------------------- So you have these two equations 3a + 2b + c = 0 -a - 3b + c = 0 it's still possible to solve this system, but it will have more than one solution

OpenStudy (anonymous):

\[ \|\mathbf v \times \mathbf w\| = \sqrt{(\|\mathbf v\|\|\mathbf w\|)^2-(\|\mathbf v\cdot \mathbf w\|)^2} \]This is my best generalization of the magnitude of the cross product What do you need it anyway?

OpenStudy (anonymous):

@wio Ya that equation makes the most sense. Oh and it's just a homework problem

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

Basically the subtraction of the dot product is trying to work against any parallelness.